# appendix,by

### v1

@@ Introduction
The study of difference equations may be carried out inside a suitable ring of functions, say in the indeterminate $x$, equipped with the endomorphism $\sigma(f(x))=f(x+1)$. For this reason, a field $K$ with a distinguished endomorphism $\sigma$ is called a difference field [Coh65]; when $\sigma$ is an automorphism, $(K,\sigma)$ is said to be inversive. The typical object of interest in this work will be an inversive valued difference field: a valued field $(K,v)$ together with an automorphism $\sigma$ compatible with the valuation $v$, that is, fixing the valuation ring $\mathcal{O}$ setwise. If we denote by $\Gamma$ the value group and by $v\colon K\to\Gamma\cup\{\infty\}$ the (surjective) valuation, we could equivalently specify a pair of automorphisms, one of the field $K$ and one of the ordered abelian group $\Gamma$, that we will both denote by $\sigma$, such that for all $a\in K$ one has $v(\sigma(a))=\sigma(v(a))$.
Automorphisms of valued fields appear for instance in [Sen69, Put72, Kel86, KS22, BG25, KKS25]. Their model theory has been studied intensively since the beginning of the millennium, for example in [Sca00, Sca07, BMS07, Azg10, AD11, Pal12, DO15, Rid17], and it has been used to obtain Lang–Weil-type estimates in finite difference fields [Hil+24] and a version of Kapranov’s Theorem for tropical difference algebra [Ali25]. A connection with algebraic dynamics was envisaged in the introduction to [CH08]. Although the techniques have recently been extended to the non-inversive case [DH22, DH25, Ram24], all difference fields in this paper will be inversive, hence we will from now on omit this qualifier and tacitly assume the surjectivity of all of our field endomorphisms, unless otherwise specified.
@@
For categories of structures with an automorphism it is customary to refer to the automorphism of an existentially closed object as generic [Mac97, CP98]. Note that not every structure carries a generic automorphism; for example, every field that does must be algebraically closed; this follows from the fact that every automorphism of a field can be extended to its algebraic closure. In the valued difference case, some consequences of genericity can be easily deduced using standard results from the model theory of valued fields: for instance (Proposition˜2.3) the valued field needs to be algebraically closed, and its residue field and value group, equipped with the induced automorphisms, need to be existentially closed in their respective categories.
@@
After observing this, one is immediately confronted with the fact that, in almost all of the existing literature on fields with a valuation and an automorphism, the behaviour of $\sigma$ on $\Gamma$ is severely constrained, and certainly not generic. Namely, most of the currently available results are only valid in the multiplicative case [Pal12], two important special cases being the isometric one —where $v(\sigma(x))=v(x)$ for every $x$, that is, the induced automorphism of $\Gamma$ is the identity— and the strongly contractive one —where $v(\sigma(x))>n\cdot v(x)$ for every $x$ of positive valuation and every $n\in\mathbb{N}$. Exceptions are [Bey+20], which deals with existentially closed $(K,v,\sigma)$ where $\sigma$ is only required to preserve the field structure but not $v$ —hence there is no induced automorphism of $\Gamma$ at all— and [DO15, Rid17, Ram24], that black-box the behaviour of $\sigma$ on $\Gamma$. Another non-multiplicative example is, on the field of transseries, precomposition by a positive infinite element, studied in [Hoe06, Chapter 5] and [DO15, Section 8].
@@
Existential closedness for ordered abelian groups with an automorphism —from now on, difference oags— has recently been studied in [DM24]. Having started to grasp an understanding of existentially closed difference oags, which occur as value groups of existentially closed valued difference fields, it is natural to wonder whether one may leverage knowledge of the former to obtain results on the latter. In this paper, we will see that the answer is affirmative.
@@
The first property one may expect from a universal domain for tropical difference algebra is the existence of a cross-section $s\colon\Gamma\to K^{\times}$ of $v$ —a group morphism $\Gamma\to K^{\times}$ such that $v(s(\gamma))=\gamma$— that is $\sigma$-equivariant. In turn, such an object induces a $\sigma$-equivariant angular component, that is, a multiplicative homomorphism from $K$ to the residue field $k$ extending the residue map. The existence of these maps implies that the inclusion of $\mathbb{Z}[\sigma]$-modules $\mathcal{O}^{\times}\hookrightarrow K^{\times}$ is pure: every system of multiplicative difference equations with parameters in $\mathcal{O}^{\times}$ that has a solution in $K^{\times}$ must have one in $\mathcal{O}^{\times}$. If said equations are not allowed to mention $\sigma$, then this is well known to follow from the fact that $\Gamma$ is torsion-free. When $\sigma$ enters the picture, showing this becomes more intricate, and constitutes our first main result.
@@
We prove the above results in Section˜2 by combining logical and tropical tools and techniques, such as the Compactness Theorem, saturation, and initial varieties. We also use a version of the Fundamental Theorem of Tropical Geometry for valuations with arbitrary rank (˜1.12), which we prove in Section˜1 by a compactness argument using upper bounds from [JS18] on the minimum degree of a tropical basis. The existence of an equivariant cross-section allows to define equivariant initial forms (˜2.16), thereby providing further evidence that this framework is suitable to develop tropical difference algebra. We will also see that, by an argument due to Philip Dittmann, existentially closed valued difference fields are not existentially closed as difference fields (Corollary˜2.27).
In Section˜3 we work in the case of residual characteristic 0, and study amalgamation problems: given a base object $A$ and morphisms $f\colon A\to B$ and $g\colon A\to C$, we seek an object $D$ and morphisms making the diagram in Figure˜1 commute.
@@
It is well known that, in the category of valued fields, as well as in that of difference fields, where in both cases we allow as arrows all possible embeddings, not every amalgamation problem can be solved. For difference fields, one may solve amalgamation problems where the arrows are required to preserve certain predicates expressing the possibility to add solutions to difference equations of a specific form [Mac97]. As for valued fields, one remedy is to use $\mathrm{RV}$-structures, or equivalently [LT25] stringent valued hyperfields; to the best of our knowledge, a theory of generic automorphisms of such structures is still lacking. At any rate, by ˜3.2, even adding the aforementioned predicates to the residue field and simultaneously including $\mathrm{RV}$-sorts does not allow to solve all amalgamation problems.
For this reason, we adopt another common approach, and work in the category of valued difference fields endowed with a $\sigma$-equivariant angular component —briefly, $\mathrm{ac}$-valued difference fields— of residual characteristic 0, with embeddings required to commute with angular components. We characterise which amalgamation problems have solutions by proving an embedding property (Proposition˜3.27) that also allows us to characterise the existentially closed objects of this category in terms of genericity on the automorphisms induced on $k$ and $\Gamma$ and a generalisation of henselianity commonly used in the context of valued difference fields.
@@
An amalgamation problem of $\mathrm{ac}$-valued difference fields of residual characteristic 0 has a solution if and only if the induced amalgamation problem of difference residue fields does.
@@
Both theorems also hold if we work with cross-sections instead of angular components, and in both cases we may also add a lift of the residue field to the language. Combining ˜C with results from [DO15] we show that, if $\operatorname{char}(k)=0$ and both $(k,\sigma)$ and $(\Gamma,\sigma)$ are existentially closed, then so is their Hahn product $k((\Gamma))$ (˜3.30).
As observed in [Pil00, Ben03, BP07], even if a class of existentially closed structures is not elementary, which happens to be the case in our setting by [KS02], one can still analyse it through the lenses of model theory by using positive logic, an approach already followed in e.g. [HK21, Kam23, EKN21, DM24] that constitutes one of the main motivations for the general study of this framework, carried out for instance in [Hay19, DK22, DGK23, Kam24]. Positive logic has also been used to introduce definability patterns [Hru20, Seg25] and obtain applications to approximate subgroups [Hru22].
The relevant neostability-theoretic dividing line for this paper is the tree property of the second kind: the existence of an unbounded definable family of independent partitions. The absence of this property is a tameness condition, known as $\mathsf{NTP}_{2}$ and studied intensively in recent years, see e.g. [Che14, CH14, MOS20, Kam24]. In our final Section˜4 we take the opportunity to advance the study of $\mathsf{NTP}_{2}$ in the positive setting by generalising to arbitrary positive theories the Array Modelling Theorem (˜4.10), a Ramsey-theoretic statement established in [Kam24] for thick positive theories. Array modelling implies a number of desirable general properties such as submultiplicativity of burden (Corollary˜4.18) that, combined with ˜B, allow us to adapt the arguments of [CH14] to prove that our class enjoys the above tameness property, failing for example in existentially closed exponential fields [HK21].
@@ Acknowledgements
Proposition˜2.23 and ˜2.17 are due respectively to Philip Dittmann and Angus Matthews; we are grateful to them for allowing us to include their results here. We thank Simone Ramello and Silvain Rideau-Kikuchi for extensive and very inspiring discussions on several problems addressed in this paper. Thanks to Martin Hils, Ehud Hrushovski and Martin Ziegler for helpful comments.
@@ 1.1 Model theory
Lowercase letters such as $a,b,x,y,\ldots$ will usually denote finite (except in Section˜4) tuples of elements or of variables, e.g. $a=({a}_{0},\ldots,{a}_{n-1})$, and their length (in the previous case, $n$) will be denoted by $\lvert a\rvert$. We abuse the notation and write e.g. $a\in M$ instead of $a\in M^{\lvert a\rvert}$. Similarly, for a function $f$ defined on $M$, we write $f(a)$ for $(f(a_{0}),\ldots,f(a_{n-1}))$, etc. We refer the reader to [DM24, Section 1.2] for a quick tour of some model-theoretic terminology and facts, to [Hod93] for a more extensive treatment of the fundamentals, and to [Dri14] for the basics of valued fields and of their model theory. We nevertheless recall the following standard definition.
@@
Let $\mathcal{L}$ be a first-order language, $T$ an $\mathcal{L}$-theory axiomatised by $\forall\exists$-sentences, and $\mathcal{K}$ the category of models of $T$ with embeddings. An object $M\in\mathcal{K}$ is existentially closed iff, for every embedding $f\colon M\to N$, every tuple $a$ from $M$, and every quantifier-free formula $\varphi(x,y)$ without parameters, if $N\vDash\exists x\;\varphi(x,f(a))$ then $M\vDash\exists x\;\varphi(x,a)$
@@
The following is well known:
@@
    Every such $\mathcal{K}$ is closed under inductive limits, and this in turn ensures that every object embeds in an existentially closed one.
@@
    By enlarging $\mathcal{L}$, hence restricting the class of embeddings between models, every first-order theory can be $\forall\exists$-axiomatised. However, this process may change the notion of existentially closed model.
@@
      a sort $\Gamma$ for the value group, carrying a copy of the language $\mathcal{L}_{\mathrm{oag}}\coloneqq\{+_{\Gamma},0_{\Gamma},-_{\Gamma},<_{\Gamma}\}$ of oags and an additional constant symbol $\infty$, to be interpreted as an infinite absorbing element equal to the valuation of $0_{\mathrm{VF}}$,
@@
    We write $\mathcal{L}_{\sigma}$ (resp., $\mathcal{L}_{\mathrm{ac},\sigma},\mathcal{L}_{s,\sigma}$) for the expansion of $\mathcal{L}_{\mathrm{vf}}$ (resp., $\mathcal{L}_{\mathrm{ac}},\mathcal{L}_{s}$) by symbols $\sigma_{\mathrm{VF}},\sigma_{\mathrm{VF}}^{-1},\sigma_{\Gamma},\sigma_{\Gamma}^{-1},\sigma_{k},\sigma_{k}^{-1}$, interpreted as automorphisms of the sorts indicated in the subscripts and their inverses, and commuting with the other maps in the language (see below).
@@
Recall that $\mathsf{DOAG}$ denotes the $\mathcal{L}_{\mathrm{oag}}$-theory of nontrivial divisible ordered abelian groups. We denote by $\mathsf{ACF}$ the $\mathcal{L}_{\mathrm{field}}$-theory of algebraically closed fields, by $\mathsf{ACVF}$ the $\mathcal{L}_{\mathrm{vf}}$-theory of nontrivially valued algebraically closed fields, by $\mathsf{ACVF}_{\mathrm{ac}}$ the $\mathcal{L}_{\mathrm{ac}}$-theory of nontrivially valued algebraically closed fields equipped with an angular component map, and similarly for the other languages mentioned above. For $p,q$ prime or zero, we write $\mathsf{ACVF}_{p,q}$ for the completion of $\mathsf{ACVF}$ obtained by specifying that $\operatorname{char}(\mathrm{VF})=p$ and $\operatorname{char}(k)=q$. As customary, in the aforementioned theories of valued fields, the valuation map $v\colon\mathrm{VF}\to\Gamma$ is assumed surjective. All of these theories are easily checked to be $\forall\exists$-axiomatisable in their language, hence each of their models embeds in an existentially closed one. [^1] In fact, all of $\mathsf{DOAG}$, $\mathsf{ACF}$, $\mathsf{ACVF}$, $\mathsf{ACVF}_{\mathrm{ac}}$ are model complete, that is, all of their models are existentially closed. After adding $\sigma,\sigma^{-1}$ to the languages above, together with axioms saying that these are respectively an automorphism and its inverse, $\forall\exists$-axiomatisability is clearly preserved.
@@
Given $(k,\sigma_{k})$ and $(\Gamma,\sigma_{\Gamma})$, the Hahn field $k((\Gamma))$ has a natural valued difference field structure, with the natural valuation and $\sigma\colon\operatorname{VF}(L)\to\operatorname{VF}(L)$ given by $(\sigma(f))(\gamma)\coloneqq\sigma_{k}(f(\sigma^{-1}_{\Gamma}(\gamma)))$ or, in power series notation,

### v2

@@ Introduction
The study of difference equations may be carried out inside a suitable ring of functions, say in the indeterminate $x$, equipped with the endomorphism $\sigma(f(x))=f(x+1)$. For this reason, a field $K$ with a distinguished endomorphism $\sigma$ is called a difference field [cohnDifferenceAlgebra1965]; when $\sigma$ is an automorphism, $(K,\sigma)$ is said to be inversive. The typical object of interest in this work will be an inversive valued difference field: a valued field $(K,v)$ together with an automorphism $\sigma$ compatible with the valuation $v$, that is, fixing the valuation ring $\mathcal{O}$ setwise. If we denote by $\Gamma$ the value group and by $v\colon K\to\Gamma\cup\{\infty\}$ the (surjective) valuation, we could equivalently specify a pair of automorphisms, one of the field $K$ and one of the ordered abelian group $\Gamma$, that we will both denote by $\sigma$, such that for all $a\in K$ one has $v(\sigma(a))=\sigma(v(a))$.
Automorphisms of valued fields appear for instance in [senAutomorphismsLocalFields1969, vanderputDifferenceEquationsOverpadic1972, kellerValuedCompleteFields1986, kuhlmannAutomorphismGroupValued2022, kaplanDecomposingAutomorphismGroup2025]. Their model theory has been studied intensively since the beginning of the millennium, for example in [Sc, Sc2, belairModelTheoryFrobenius2007, Az, AvD, Pa, DO, Ri, blaszkiewiczNoteValuedFields2025], and it has been used to obtain Lang–Weil-type estimates in finite difference fields [hilsLangWeilTypeEstimates2024] and a version of Kapranov’s Theorem for tropical difference algebra [aliyariDifferenceKapranovTheorem2025]. A connection with algebraic dynamics was envisaged in the introduction to [chatzidakisDifferenceFieldsDescent2008]. Although the techniques have recently been extended to the non-inversive case [dorSpecializationDifferenceEquations2022, dorContractingEndomorphismsValued2023, ramelloModelTheoryValued2024], all difference fields in this paper will be inversive, hence we will from now on omit this qualifier and tacitly assume the surjectivity of all of our field endomorphisms, unless otherwise specified.
@@
For categories of structures with an automorphism it is customary to refer to the automorphism of an existentially closed object as generic [Ma, chatzidakisGenericStructuresSimple1998]. Note that not every structure carries a generic automorphism; for example, every field that does must be algebraically closed; this follows from the fact that every automorphism of a field can be extended to its algebraic closure. In the valued difference case, some consequences of genericity can be easily deduced using standard results from the model theory of valued fields: for instance (Proposition˜2.3) the valued field needs to be algebraically closed, and its residue field and value group, equipped with the induced automorphisms, need to be existentially closed in their respective categories.
@@
After observing this, one is immediately confronted with the fact that, in almost all of the existing literature on fields with a valuation and an automorphism, the behaviour of $\sigma$ on $\Gamma$ is severely constrained, and certainly not generic. Namely, most of the currently available results are only valid in the multiplicative case [Pa], two important special cases being the isometric one—where $v(\sigma(x))=v(x)$ for every $x$, that is, the induced automorphism of $\Gamma$ is the identity—and the strongly contractive one—where $v(\sigma(x))>n\cdot v(x)$ for every $x$ of positive valuation and every $n\in\mathbb{N}$. Exceptions are [beyarslanFieldsAutomorphismValuation2020], which deals with existentially closed $(K,v,\sigma)$ where $\sigma$ is only required to preserve the field structure but not vv—hence there is no induced automorphism of $\Gamma$ at all—and [DO, Ri, ramelloModelTheoryValued2024], that black-box the behaviour of $\sigma$ on $\Gamma$. Another non-multiplicative example is, on the field of transseries, precomposition by a positive infinite element, studied in [vanderhoevenTransseriesRealDifferential2006a, Chapter 5] and [DO, Section 8].
@@
Existential closedness for ordered abelian groups with an automorphism—from now on, difference oags—has recently been studied in [DM]. Having started to grasp an understanding of existentially closed difference oags, which occur as value groups of existentially closed valued difference fields, it is natural to wonder whether one may leverage knowledge of the former to obtain results on the latter. In this paper, we will see that the answer is affirmative.
@@
The first property one may expect from a universal domain for tropical difference algebra is the existence of a cross-section of vv—a group morphism $s\colon\Gamma\to K^{\times}$ such that $v(s(\gamma))=\gamma$—that is $\sigma$-equivariant. In turn, such an object induces a $\sigma$-equivariant angular component, that is, a multiplicative homomorphism from $K$ to the residue field $k$ extending the residue map. The existence of these maps implies that the inclusion of $\mathbb{Z}[\sigma]$-modules $\mathcal{O}^{\times}\hookrightarrow K^{\times}$ is pure: every system of multiplicative difference equations with parameters in $\mathcal{O}^{\times}$ that has a solution in $K^{\times}$ must have one in $\mathcal{O}^{\times}$. If said equations are not allowed to mention $\sigma$, then this is well known to follow from the fact that $\Gamma$ is torsion-free. When $\sigma$ enters the picture, showing this becomes more intricate, and constitutes our first main result.
@@
We prove the above results in Section˜2 by combining logical and tropical tools, such as the Compactness Theorem, saturation, and initial varieties. We also use a version of the Fundamental Theorem of Tropical Geometry for valuations with arbitrary rank (˜1.12), which we prove in Section˜1 by a compactness argument using upper bounds from [joswigDegreeTropicalBasis2018] on the minimum degree of a tropical basis. The existence of an equivariant cross-section allows to define equivariant initial forms (˜2.16), thereby providing further evidence that this framework is suitable to develop tropical difference algebra. We will also see that, by an argument due to Philip Dittmann, existentially closed valued difference fields are not existentially closed as difference fields (Corollary˜2.27).
In Section˜3 we work in the case of residual characteristic zero, and study amalgamation problems: given a base object $A$ and morphisms $f\colon A\to B$ and $g\colon A\to C$, we seek an object $D$ and morphisms making the diagram in Figure˜1 commute.
@@
It is well known that, in the category of valued fields, as well as in that of difference fields, where in both cases we allow as arrows all possible embeddings, not every amalgamation problem can be solved. For difference fields, one may solve amalgamation problems where the arrows are required to preserve certain predicates expressing the possibility to add solutions to difference equations of a specific form [Ma]. In our setting, even adding the aforementioned predicates to the residue field does not allow to solve all amalgamation problems (by ˜3.2).
As for valued fields, one remedy is to use $\mathrm{RV}$-structures, or equivalently [linziHyperfieldsAssociatedValued2025] stringent valued hyperfields; to the best of our knowledge, a theory of generic automorphisms of such structures is still lacking. For this reason, we adopt another common approach, and work in the category of valued difference fields endowed with a $\sigma$-equivariant angular component—briefly, $\mathrm{ac}$-valued difference fields—of residual characteristic zero, with embeddings required to commute with angular components. We characterise which amalgamation problems have solutions by proving an embedding result (Proposition˜3.27) that also allows us to characterise the existentially closed objects of this category in terms of genericity of the automorphisms induced on $k$ and $\Gamma$ and a generalisation of henselianity commonly used in the context of valued difference fields.
@@
An amalgamation problem of $\mathrm{ac}$-valued difference fields of residual characteristic zero has a solution if and only if the induced amalgamation problem of difference residue fields does.
@@
Both theorems also hold if we work with cross-sections instead of angular components, and in both cases we may also add a lift of the residue field to the language. Combining ˜C with results from [DO] we show that, if $\operatorname{char}(k)=0$ and both $(k,\sigma)$ and $(\Gamma,\sigma)$ are existentially closed, then so is their Hahn product $k((\Gamma))$ (˜3.30).
As observed in [pillay_forking_2000, BY, BYP], even if a class of existentially closed structures is not elementary, which happens to be the case in our setting by [kikyoStrictOrderProperty2002], one can still analyse it through the lenses of model theory by using positive logic, an approach already followed in e.g. [HK, kamsmaBilinearSpacesFixed2023, DKN, DM] that constitutes one of the main motivations for the general study of this framework, carried out for instance in [Ha, DK, DGK, Ka]. Positive logic has also been used to introduce definability patterns [hrushovskiDefinabilityPatternsTheir2020, segelPositiveDefinabilityPatterns2025] and obtain applications to approximate subgroups [hrushovskiLascarGroup2022].
The relevant neostability-theoretic dividing line for this paper is the tree property of the second kind: the existence of an unbounded definable family of independent partitions. The absence of this property is a tameness condition, known as $\mathsf{NTP}_{2}$ and studied intensively in recent years, see e.g. [Chernikov, CH, montenegroSTABILIZERSGROUPSGENERICS2020, Ka]. In our final Section˜4 we take the opportunity to advance the study of $\mathsf{NTP}_{2}$ in the positive setting by generalising to arbitrary positive theories the Array Modelling Theorem (˜4.10), a Ramsey-theoretic statement established in [Ka] for thick positive theories. Array modelling implies a number of desirable general properties such as submultiplicativity of burden (Corollary˜4.18) that, combined with ˜B, allow us to adapt the arguments of [CH] to prove that our class enjoys the above tameness property, which for example fails in existentially closed exponential fields [HK].
@@ Acknowledgements
Proposition˜2.23 and ˜2.17 are due respectively to Philip Dittmann and Angus Matthews; we are grateful to them for allowing us to include their results here. We thank Simone Ramello and Silvain Rideau-Kikuchi for extensive and very inspiring discussions on several problems addressed in this paper. Thanks to Philip Dittmann, Martin Hils, Ehud Hrushovski, Mark Kamsma and Martin Ziegler for helpful comments.
@@ 1.1 Model theory
Lowercase letters such as $a,b,x,y,\ldots$ will usually denote finite (except in Section˜4) tuples of elements or of variables, e.g. $a=({a}_{0},\ldots,{a}_{n-1})$, and their length (in the previous case, $n$) will be denoted by $\lvert a\rvert$. We abuse the notation and write e.g. $a\in M$ instead of $a\in M^{\lvert a\rvert}$. Similarly, for a function $f$ defined on $M$, we write $f(a)$ for $(f(a_{0}),\ldots,f(a_{n-1}))$, etc. We refer the reader to [DM, Section 1.2] for a quick tour of some model-theoretic terminology and facts, to [Ho] for a more extensive treatment of the fundamentals, and to [Dri] for the basics of valued fields and of their model theory. We nevertheless recall the following standard definition.
@@
Let $\mathcal{L}$ be a first-order language, $T$ an $\mathcal{L}$-theory axiomatised by $\forall\exists$-sentences, and $\mathcal{K}$ the category of models of $T$ with embeddings. An object $M\in\mathcal{K}$ is existentially closed iff, for every embedding $f\colon M\to N$, every tuple $a$ from $M$, and every quantifier-free formula $\varphi(x,y)$ without parameters, if $N\vDash\exists x\;\varphi(x,f(a))$ then $M\vDash\exists x\;\varphi(x,a)$.
@@
The following is well known.
@@
    Every such $\mathcal{K}$ is closed under inductive limits [Ho, Theorem 6.5.9 and Exercise 5 in Section 6.5], and this in turn ensures that every object embeds in an existentially closed one [Ho, Theorem 8.2.3].
@@
    By enlarging $\mathcal{L}$, hence restricting the class of embeddings between models, every first-order theory can be $\forall\exists$-axiomatised [Ho, Theorems 2.6.5 and 2.6.6]. However, this process (known as Morleyisation or atomisation) may change the notion of existentially closed model.
@@
      a sort $\Gamma$ for the value group, carrying a copy of the language $\mathcal{L}_{\mathrm{oag}}\coloneqq\{+_{\Gamma},0_{\Gamma},-_{\Gamma},<_{\Gamma}\}$ of oags and an additional constant symbol $\infty$, to be interpreted as an infinite absorbing element, with $\infty=\infty$, equal to the valuation of $0_{\mathrm{VF}}$,
@@
    We write $\mathcal{L}_{\sigma}$ (resp. $\mathcal{L}_{\mathrm{ac},\sigma},\mathcal{L}_{s,\sigma}$) for the expansion of $\mathcal{L}_{\mathrm{vf}}$ (resp. $\mathcal{L}_{\mathrm{ac}},\mathcal{L}_{s}$) by symbols $\sigma_{\mathrm{VF}},\sigma_{\mathrm{VF}}^{-1},\sigma_{\Gamma},\sigma_{\Gamma}^{-1},\sigma_{k},\sigma_{k}^{-1}$, interpreted as automorphisms of the sorts indicated in the subscripts and their inverses, and commuting with the other maps in the language (see below).
@@
Recall that $\mathsf{DOAG}$ denotes the $\mathcal{L}_{\mathrm{oag}}$-theory of nontrivial divisible ordered abelian groups. We denote by $\mathsf{ACF}$ the $\mathcal{L}_{\mathrm{field}}$-theory of algebraically closed fields, by $\mathsf{ACVF}$ the $\mathcal{L}_{\mathrm{vf}}$-theory of nontrivially valued algebraically closed fields, by $\mathsf{ACVF}_{\mathrm{ac}}$ the $\mathcal{L}_{\mathrm{ac}}$-theory of nontrivially valued algebraically closed fields equipped with an angular component map, and similarly for the other languages mentioned above. For $p,q$ prime or zero, we write $\mathsf{ACVF}_{p,q}$ for the completion of $\mathsf{ACVF}$ obtained by specifying that $\operatorname{char}(\mathrm{VF})=p$ and $\operatorname{char}(k)=q$. As customary, in the aforementioned theories of valued fields, the valuation map $v\colon\mathrm{VF}\to\Gamma$ is assumed surjective. All of these theories are easily checked to be $\forall\exists$-axiomatisable in their language, hence by ˜1.2 each of their models embeds in an existentially closed one. [^1] In fact, all of $\mathsf{DOAG}$, $\mathsf{ACF}$, $\mathsf{ACVF}$, $\mathsf{ACVF}_{\mathrm{ac}}$ are model complete, that is, all of their models are existentially closed. After adding $\sigma,\sigma^{-1}$ to the languages above, together with axioms saying that these are respectively an automorphism and its inverse, $\forall\exists$-axiomatisability is clearly preserved.
@@
Given $(k,\sigma_{k})$ and $(\Gamma,\sigma_{\Gamma})$, the Hahn field $k((\Gamma))$ can be equipped with the structure of a valued difference field, with the natural valuation and $\sigma\colon\operatorname{VF}(L)\to\operatorname{VF}(L)$ given by $(\sigma(f))(\gamma)\coloneqq\sigma_{k}(f(\sigma^{-1}_{\Gamma}(\gamma)))$ or, in power series notation,